OpenStudy (anonymous):

If , f(x)= 2x^2 +5 times the square root of (x-2) complete the following statement: f(3) = _____

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

\(\ \sf f(x) = 2x^2 + 5 \times \sqrt{x - 2} \) ?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup i got 23 too

OpenStudy (the_fizicx99):

kk, f(3) means to substitute 3 everywhere you see an "x" so, \(\ \sf 2(3)^2 + 5 \times \sqrt{3 - 2} = 2*9 + 5 * 1 = 18 + 5 * 1 = 23 * 1 = 23 \)

OpenStudy (anonymous):

thank you!

OpenStudy (the_fizicx99):

Yw

OpenStudy (anonymous):

would anyone mind taking a look at this one?

OpenStudy (the_fizicx99):

A and D

OpenStudy (anonymous):

Thanks a bunch!